OpenStudy (purple_pink):

find 2 ratios that are equivalent to the given ratio 6 : 10

OpenStudy (maddisongruby):

3:5 60:100 12:20 18:30

OpenStudy (theprofessor35):

Lets multiply!

OpenStudy (theprofessor35):

6*2 = 12 10*2 = 20

OpenStudy (theprofessor35):

Now, what about you try another?

OpenStudy (theprofessor35):

6*3 = ? 10*3 = ?

OpenStudy (purple_pink):

where'd you get the 2?

OpenStudy (theprofessor35):

You multiply by ANY number, as long as you do it to BOTH.

OpenStudy (theprofessor35):

Like this: 6:10 = 12:20 U know why? 6*2 = 12 10*2 = 20 12:20

OpenStudy (purple_pink):

ohhh

OpenStudy (theprofessor35):

Get it now?

OpenStudy (purple_pink):

yep thanks

OpenStudy (theprofessor35):

:) Yay!